solve the equation. if needed 3 decimal places

77=100-100e^-0.09x
Answer #1

Solution. The given eqo is -0.0px 77 = 100 - 100 e -0.oga -> 100e = 23 = -0.ogx e = 0.23 100 = -0.09x = ls (0.23) =) - r = ln